# Is there a dynamical tendency in $H_{0}$ with late time measurements?

### v1

@@ 6 Summary and conclusions
The SNe observations are chosen from one of the three most recent samples of Pantheon+, Union3 and DES. Pantheon+ and Union 3 share about 1360 SNe but differ in their analysis and treatment over systematics. DES has 1635 new SNe but share 194 low redshift SNe with Pantheon+. Our three data samples share the megamasers+CC+DESI data but differ in the SNe data. Base+PantheonPlus and Base+Union3 formed a total of 8 bins, and 6 bins for Base+DES.
@@
The results show a dynamical trend in $H_{0}$ with a statistical significance up to $1.5\sigma$, for the quadratic parameterization (5.6) and up to $2.3\sigma$ for the Fourier parameterization (5.8), indicating a preference for a non constant model. However, the values in $\Delta$AIC and the Bayes factor $B$, which penalize the extra parameters in the dynamical models, do not provide substantial support for the quadratic function against a constant value. From a Bayesian point of view, both the quadratic function and a constant $H_{0}$ fit the bins results equally well. Similar conclusions are obtained for the Fourier parameterization, except for the results of Base+DES, which have the highest $\Delta$AIC and $B$ values showing a substantial preference for the dynamical model, according to the Jeffrey’s scale.
@@
We highlight that the three data sets, which use different SNe samples, show similar behaviors even though each sample differs in its content and methodology. Because of this, we can infer that this trend is not caused by uncountable systematic errors in the SNe parameters and it is more of an intrinsic behavior guided by the data.
@@
We have ignored any calibration on either $r_{d}$ or M. However, if we adopt them, we find that the bins values of $H_{0}$ are completely consistent with a constant. This would require to use data outside of the redshift bin, either CMB data to determine $r_{d}$ or distance leader data to determine M. As fixing $r_{d}$ or M is equivalent to fixing $H_{0}$ from the beginning, which prevents the data from speaking for itself. It is relevant to highlight that any solution to the Hubble tension that only modifies the calibration of $r_{d}$ or M won’t solve the tension presented here. In the unbinned case, in subsection 5.3 we obtained a $r_{d}$ which is $1.5\sigma$ smaller than the one measured by Planck.
@@
Of course, the results for $H_{0}$ may vary if one adopts a different binning strategy but we expect that the significance of the trend do not change significantly. This evolution in the Hubble constant might offer a new perspective on the Hubble tension. Since we’re using the $w_{0}w_{a}$ model with very broad priors, it seems that a dynamical DE can’t remove this trend. Using a cosmological constant as the Dark Energy component only reduces the errorbars of the estimation of $H_{0}$ due to the reduced parameter space. This may increase the significance of the non constant models.
@@
Some possible solutions would be either a hidden systematic error or some other assumption in the model. This could even be the reason why SH0ES and Planck have different values for $H_{0}$ given that they use data with different redshifts, but this needs further study, maybe considering $H_{0}$ not as an integration constant, but as a quantity dependent on the observed region of the Universe. Unfortunately, directly corroborating that the Hubble parameter at present time has different values depending on the region of the Universe is difficult, as due to the finiteness of the speed of light, we observe those regions in the past. However, nothing prevents us from thinking that just as there are perturbations to matter, these bring with them anisotropies in the expansion rate of the Universe with $H_{0}$ being more of an average value and the local departure of the cosmological principle [65, 66, 67, 68, 69, 70].
@@
Our binning strategy splits the data used, reducing the available data in each bin. This, combined with the decreasing quality in the data, produces large error bars in the inferred parameters which get worse at large $z$’s. Therefore the quadratic and Fourier fittings present large uncertainties. We do not rule out that the constant $H_{0}$ be the correct result; we need to wait to improve the quality and number of the data to be able to reach a final conclusion.
@@
The Hubble tension remains one of the most significant challenges in modern cosmology and astrophysics, highlighting persistent discrepancies in $H_{0}$ values derived from different observational methods and across varied redshifts. This work contributes to the present discussion, providing further evidence that the Hubble tension is not only confined to early- versus late-time measurements but also appears consistently within purely late-time measurements reflecting unresolved aspect of cosmological models and data calibration.
@@ Acknowledgments
The authors thankfully acknowledge Sebastien Fromenteau for his useful comments in the development of this work. M. L. H. was supported by Conahcyt grant 806098.

### v2

@@ 6 Summary and conclusions
SNe observations are chosen from one of the three most recent samples of Pantheon+, Union3, and DES. Pantheon+ and Union 3 share about 1360 SNe, but differ in their analysis and treatment over systematics. DES has 1635 new SNe but shares 194 low-redshift SNe with Pantheon+. Our three data samples share the megamasers+CC+DESI data, but differ in the SNe data. Base+PantheonPlus and Base+Union3 formed a total of 8 bins, and 6 bins for Base+DES.
@@
The results show a dynamical trend in $H_{0}$ with a statistical significance up to $1.5\sigma$, for the quadratic parameterization (5.6) and up to $2.3\sigma$ for the Fourier parameterization (5.8), indicating a preference for a non-constant model. However, the values in $\Delta$AIC and the Bayes factor $B$, which penalize the extra parameters in the dynamical models, do not provide substantial support for the quadratic function against a constant value. From a Bayesian point of view, both the quadratic function and a constant $H_{0}$ fit the bins results equally well. Similar conclusions are obtained for the Fourier parameterization, except for the results of Base+DES, which have the highest $\Delta$AIC and $B$ values showing a substantial preference for the dynamical model, according to Jeffrey’s scale.
@@
We highlight that the three datasets, which use different SNe samples, show similar behaviors even though each sample differs in its content and methodology. Because of this, we can infer that this trend is not caused by uncountable systematic errors in the SNe parameters, and it is more an intrinsic behavior guided by the data.
@@
We have ignored any calibration on either $r_{d}$ or M. However, if we adopt them, we find that the bins values of $H_{0}$ are completely consistent with a constant. This would require using data outside of the redshift bin, either CMB data to determine $r_{d}$ or distance leader data to determine M. As fixing $r_{d}$ or M is equivalent to fixing $H_{0}$ from the beginning, which prevents the data from speaking for itself. It is relevant to highlight that any solution to the Hubble tension that only modifies the calibration of $r_{d}$ or M will not solve the tension presented here. In the unbinned case, in subsection 5.3 we obtained a $r_{d}$ that is $1.5\sigma$ smaller than the one measured by Planck.
@@
Of course, the results for $H_{0}$ may vary if one adopts a different binning strategy, but we expect that the significance of the trend does not change significantly. This evolution in the Hubble constant might offer a new perspective on the Hubble tension. Since we are using the $w_{0}w_{a}$ model with very broad priors, it seems that a dynamical DE cannot remove this trend. It’s important to note that using a cosmological constant as the Dark Energy component only reduces the errorbars in the estimation of $H_{0}$ due to the reduced parameter space. This may increase the significance of the non-constant models.
@@
Some possible solutions would be either a hidden systematic error or some other assumption in the model. This could even be the reason why SH0ES and Planck have different values for $H_{0}$ given that they use data with different redshifts, but this needs further study, maybe considering $H_{0}$ not as an integration constant, but as a quantity dependent on the observed region of the Universe. Unfortunately, directly corroborating that the Hubble parameter at the present time has different values depending on the region of the Universe is difficult; due to the finiteness of the speed of light, we observe those regions in the past. However, nothing prevents us from thinking that just as there are perturbations to matter, these bring with them anisotropies in the expansion rate of the Universe, with $H_{0}$ being more of an average value and the local departure of the cosmological principle [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90]. See [91] for a review that surveys observational hints for deviations of the cosmological principle.
We performed a reconstruction of the Hubble parameter, $H(z)$, as an alternative strategy to binning the data, treating it as a free function determined by its values at specific redshift nodes, $H(z_{i})$, which were used as fitting parameters. We then interpolated these values using cubic splines to obtain a continuous function for $H(z)$. Our analysis of the reconstructed Hubble parameter reveals deviations from the theoretical prediction under the CPL model, with Base+Union3 showing the most significant preference for the reconstructed $H(z)$, reaching a statistical significance of up to $3.5\sigma$ and a $\Delta$AIC value that indicates considerable support. Additionally, a slight tension is observed around $z\sim 0.5$ in Base+PantheonPlus and Base+DES between the reconstructed model and the CPL predictions. While checking different interpolation methods, including quadratic and linear splines, did not significantly our results.
@@
Our binning strategy splits the data used, reducing the available data in each bin. This, combined with the decreasing quality in the data, produces large error bars in the inferred parameters which get worse at large $z$’s. Therefore, the quadratic and Fourier fittings present large uncertainties. We do not rule out that the constant $H_{0}$ is the correct result; we need to wait to improve the quality and number of the data to be able to reach a final conclusion.
@@
The Hubble tension remains one of the most significant challenges in modern cosmology and astrophysics, highlighting persistent discrepancies in $H_{0}$ values derived from different observational methods and across various redshifts. This work contributes to the present discussion, providing further evidence that the Hubble tension is not only confined to early- versus late-time measurements but also appears consistently within purely late-time measurements reflecting unresolved aspect of cosmological models and data calibration.
@@ Acknowledgments
The authors thankfully acknowledge Sebastien Fromenteau for his useful comments in the development of this work. M. L. H. was supported by Secihti grant 806098.